OpenStudy (anonymous):

suppose cos(s) = (-1/5) and sin(t) = (3/5). Suppose also that s is in quadrant 2 and t is in quadrant 2. what is: sin(s+t) ? cos(s+t) ? tan(s+t) ?

OpenStudy (anonymous):

first of all you need cost and sin s. do you know how to get it? then you have formula for those, just apply, sin (s+t) = sins cos t+ sint cos s

OpenStudy (anonymous):

i dont know how to get the cos t or sin s :(

OpenStudy (anonymous):



OpenStudy (anonymous):

sin t = opp/ hyp = 1/5 it means you have hp =5 and opp =1 . use pythagorean to find adj .

OpenStudy (anonymous):

when having adj, cos t = adj/ hyp you got cost. the same way to find sin s . and then replace all